OpenStudy (anonymous):

How do you set up this equation? Ann's age is seven tenths of Bill's age. In 4 years Ann's age will be eight-elevenths of Bill's age. How old is Bill now.

OpenStudy (anonymous):

x = bill's age 7x/10 = ann's age

OpenStudy (anonymous):

Bill's age = x Anne's age = (7/10)x 4 years later (both have aged 4 years): Anne's age + 4 = (8/11)(Bill's age + 4) (7/10)x + 4 = (8/11)(x + 4) x = 40